OpenStudy (jojokiw3):

If (x-4)^2 + 4(y-3)^2 = 16, find the sum of distances from any fixed point on the curve of two foci

OpenStudy (jackthegreatest):

I believe you divide both sides by 16

OpenStudy (jojokiw3):

So this? \[\frac{ (x-4)^2 }{ 16 } + \frac{ (y-3)^2 }{ 4 } = 1\]

OpenStudy (jackthegreatest):

yep, then you find the horizontal and vertical axis, (a^2=16), (b^2=4) ---->

OpenStudy (jojokiw3):

This is the step that I get stuck on.

OpenStudy (jackthegreatest):

ahh i get it now

OpenStudy (jackthegreatest):

16 is the major radius/axis, the square root is 4 ( to the center), so you multiply it by 2 and you get 8

OpenStudy (jackthegreatest):

@jojokiw3

OpenStudy (jackthegreatest):

do you kind of get it now?